# PSR J1922+37: a 1.9-second pulsar discovered in the direction of the old open cluster NGC 6791

### v1

@@ 5 Discussion
The small angular separation and the close distance predicted by the NE2001 model make PSR J1922+37 a likely member of NGC 6791. Actually, three Globular cluster pulsars are significantly far from corresponding cluster center, including PSRs B1718--19A (NGC 6342, Lyne et al., 1993), J1911--5958A (NGC 6752, D’Amico et al., 2001) and J1801--0857D (NGC 6517, Lynch et al., 2011). A fourth pulsar, J1823--3022, is at $\sim 3$ times the half-light radius of NGC 6624, although their association still awaits confirmation (Abbate et al., 2022).
If the association is confirmed, PSR J1922+37 will be the first pulsar found in an open cluster, and more pulsars may be discovered in NGC 6791. However, the large discrepancy between $d_{\rm YMW16}$ and the distance to NGC 6791 suggests that a coincidental overlap is also possible. Considering the large uncertainty of distance estimates from electron density models, it is vital to check the association with further investigations.
There are several ways to check if PSR J1922+37 is actually associated with NGC 6791. First, a precise measurement of the distance to the pulsar and the open cluster will give a straightforward answer to the problem. The recent Gaia data release has allowed a much better estimation of the distance to NGC 6791 (e.g. Dias et al., 2021; Hunt & Reffert, 2024), and the estimate will likely improve in future data releases. However, the low flux density of PSR J1922+37 will pose a challenge to measuring its distance with interferometry. The large distance also makes measurement with timing parallax difficult.
Second, if PSR J1922+37 is associated with NGC 6791, then both objects should have similar proper motion. Bedin et al. (2006) reported a proper motion of $(\mu_{\alpha}\cos\delta,\mu_{\delta})=(-0.57,-2.45)$ mas yr-1 for NGC 6791, while Hunt & Reffert (2024) gave $(-0.42,-2.28)$ mas yr-1. A measurement of the proper motion of PSR J1922+37, e.g. via pulsar timing, will be helpful in determining the relation between the two objects.
@@ 5.1 A pulsar population in NGC 6791?
NGC 6791 is an unusually old and massive open cluster in our Galaxy (Grundahl et al., 2008), with an age of several Gyr (Dias et al., 2021; Hunt & Reffert, 2024) and a total mass ranging from 5000 M (Platais et al., 2011) to a few $10^{4}$ M (Hunt & Reffert, 2024). The cluster is relatively compact and follows a single-mass King density profile up to $\sim 600\arcsec$, with a core region of $r_{\rm c}\sim 2\farcm 66$, compactness of $c\sim 1.06$ and a central surface brightness of $\mu\sim 17.5$ mag per asrcsec2 (Dalessandro et al., 2015). The relatively high surface brightness and compactness thus lead to a crowded stellar environment, making NGC 6791 a possible birthplace of pulsars, similar to the pulsars in globular clusters (Verbunt, 2002; Hui et al., 2010; Bahramian et al., 2013).
@@
Following Bahramian et al. (2013), we estimate the expected number of pulsar discoveries, $\lambda$, in NGC 6791 in three steps. First, using the cluster parameters above, we computed the central luminosity density, $\rho_{\rm c}$, according to the prescription given by Djorgovski (1993). Second, we obtained the encounter rate $\Gamma\sim\rho_{\rm c}^{2}r_{\rm c}^{3}/\sigma_{v}\approx 162$ assuming a King potential model (Dalessandro et al., 2015) and a velocity dispersion of $\sigma_{v}\sim 2$ km s-1. Note that $\Gamma$ has been normalized so that the encounter rate of 47 Tuc is $\Gamma_{\rm 47~{}Tuc}=1000$ (Bahramian et al., 2013). Finally, the expected number of pulsars is found by $\ln\lambda=-1.1+1.5\log_{10}\Gamma$ (Turk & Lorimer, 2013; Gao et al., 2024), which gives $\lambda\sim 9$ potential discoveries. This result is also in general agreement with the estimation using escape velocity; see Yin et al. (2024, Fig. 2). Therefore, more pulsars are likely in NGC 6791, and it deserves future observations.
5.2 Potential pulsar discoveries in other open clusters
@@
The possible association motivates us to check if there are other open clusters that may also host pulsars. A detailed analysis inspecting the association between all known open clusters and pulsars is beyond the scope of this work, and will be presented in Zhou et al. (2024, in preparation). Here, we check the latest Galactic cluster catalog (Hunt & Reffert, 2024) and highlight the open clusters that are worth a careful pulsar search.
@@
We cannot obtain $\lambda$ for each open cluster due to the unavailability of $\rho_{\rm c}$. Instead, we sift through the catalog and pick out the open clusters that are similar to NGC 6791 in both total mass ($>10^{4}$ M) and core size ($<10$ pc), selecting open clusters with comparable core density and thus encounter rate. We also set a lower age limit of 100 Myr, which is a typical lifetime for canonical pulsars.
@@
Ten open clusters, including NGC 6791, meet the above conditions and are shown in Table 4, which lists the total mass, radial velocity scattering, angular size, and number of member stars. All the clusters are relatively close ($d<5$ kpc) and have a small core size with a typical value of $\sim 5\arcmin$. They also have a relatively high star concentration, with half of members within the tidal radius. These clusters are thus good candidates for future pulsar search in open clusters.
@@
Table 4: Open clusters similar to NGC 6791 in decreasing total mass. Columns are cluster name, right ascention and declination, distance, age, total mass, the standard error of radial velocity, core radius (in pc and in arcmin), tidal radius, total radius, number of member stars within $r_{\rm tidal}$ and the total number of stars.
6 Conclusions
Pulsars located or born in open clusters have not drawn much attention in pulsar research, but may hold important clues in understanding the formation and evolution of different species of pulsars and star clusters. In this paper, we present initial results from 20-hour pulsar searching observations in seven open clusters using FAST. We discovered a new pulsar J1922+37 in the direction of open cluster NGC 6791 (Fig. 2); it has a period of 1.92 second and a DM of $\sim 85$ pc cm-3. The NE2001 model gives a distance estimate of 4.79 kpc to the pulsar, consistent with the cluster distance, which ranges from 4.02 to 4.93 kpc in the literature. The pulsar is thus a likely member of NGC 6791. If confirmed, PSR J1922+37 will be the first pulsar discovered in an open cluster.
Measurements of pulsar distance and proper motion will be crucial in determining whether PSR J1922+37 is associated with NGC 6791 or not. Assuming that old and massive open clusters such as NGC 6791 are similar to Galactic globular clusters as pulsar birth places (e.g., the size of pulsar population is proportional to the stellar encounter rate), we estimate that around 9 pulsars are expected to be discovered in NGC 6791. We also highlight 10 open clusters that are promising sites for finding new pulsars.
Follow-up observations have been planned with FAST, which will allow measurements of spin and astrometric parameters of PSR J1922+37 as well as a deeper search for more pulsars in NGC 6791. More pulsar discoveries with similar DM to PSR J1922+37 in the region will provide compelling evidence for pulsars in open clusters.
We thank Lei Zhang, Pei Wang, Zhichen Pan, Lei Qian and Zucheng Chen for helpful discussions. This work made use of the data from FAST (Five-hundred-meter Aperture Spherical radio Telescope) (https://cstr.cn/31116.02.FAST). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. The data processing was mainly performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government. R.P.E. is supported by the Chinese Academy of Sciences President’s International Fellowship Initiative, grant No. 2021FSM0004. JPY, NW, WWZ are supported by the National Natural Science Foundation of China (grant No. 12041304, 12041303). WWZ is supported by National SKA Program of China (grant No. 2020SKA0120200). XJZ is supported by the National Natural Science Foundation of China (Grant No. 12203004) and by the Fundamental Research Funds for the Central Universities.

### v2

@@ 5 Discussion
The close vlaue between $d_{\rm NE2001}$ and $d_{\rm NGC6791}$ makes PSR J1922+37 a plausible member at the edge of the NGC 6791. In fact, there are three Globular cluster pulsars that reside reasonably far from their corresponding cluster centers, including PSRs B1718--19A (NGC 6342, Lyne et al., 1993), J1911--5958A (NGC 6752, D’Amico et al., 2001) and J1801--0857D (NGC 6517, Lynch et al., 2011). A fourth pulsar, J1823--3022, is at $\sim 3$ times the half-light radius of NGC 6624, although its association still awaits confirmation (Abbate et al., 2022).
If our open cluster association is confirmed, PSR J1922+37 will be the first radio pulsar found in an open cluster, and more may be discovered in NGC 6791. However, the large discrepancy between $d_{\rm YMW16}$ and $d_{\rm NGC6791}$ suggests that a coincidental spatial overlap is also a possibility. Considering the large uncertainty of distance estimates from electron density models, it is vital to check the association with further evidence.
There are several ways to check if PSR J1922+37 is actually associated with NGC 6791. First, a precise measurement of the distance to the pulsar and the open cluster will give a straightforward answer to the problem. The recent Gaia data release has allowed a much better estimation of the distance to NGC 6791 (e.g. Dias et al., 2021; Hunt & Reffert, 2024), and the estimate will likely improve in future data releases. However, the low flux density of PSR J1922+37 will pose a challenge in measuring its distance with long baseline interferometry. No pulsar with flux density $<10~{}\mu$Jy has ever had its parallax determined by interferometry. Long-period pulsars also exhibit significant timing noise that makes timing parallaxes impossible, so without a long baseline array that includes FAST it is difficult to see the pulsar’s distance being measurable in the near future. A distance constraint through HI absorption measurements might be possible.
Secondly, if PSR J1922+37 is associated with NGC 6791, then both objects should have similar proper motions. Bedin et al. (2006) reported a proper motion of $(\mu_{\alpha}\cos\delta,\mu_{\delta})=(-0.57,-2.45)$ mas yr-1 for NGC 6791, while Hunt & Reffert (2024) gave $(-0.42,-2.28)$ mas yr-1. A measurement of the proper motion of PSR J1922+37, e.g. via pulsar timing, would be helpful in determining the physical relation between the two objects, but timing proper motions on long-period pulsars are extremely difficult on anything except decadal timescales.
@@ 5.1 Formation mechanism
NGC 6791 is an unusually old and massive open cluster in our Galaxy (Grundahl et al., 2008), with an age of several Gyr (Dias et al., 2021; Hunt & Reffert, 2024) and a total mass ranging from 5000 M (Platais et al., 2011) to a few $10^{4}$ M (Hunt & Reffert, 2024). The cluster is relatively compact and follows a single-mass King density profile up to $\sim 600\arcsec$, with a core region of $r_{\rm c}\sim 2\farcm 66$, compactness of $c\sim 1.06$ and a central surface brightness of $\mu\sim 17.5$ mag per arcsec2 (Dalessandro et al., 2015). The relatively high surface brightness and compactness thus lead to a crowded stellar environment, making NGC 6791 a possible birthplace of pulsars, similar to globular clusters (Verbunt, 2002; Hui et al., 2010; Bahramian et al., 2013). If J1922+37 is associated with NGC 6791, it then becomes interesting to understand its formation mechanism.
If the pulsar is a typical slow pulsar, then in order to be associated with the cluster the pulsar must have been born in the initial wave of star formation many Gyr ago, have been retained by the cluster despite any natal kick, and still be an active pulsar, and be beamed at the Sun. This is improbable. $10^{4}$ M clusters only produce of order 50 neutron stars, and due to the low escape velocities (a few km s-1) most will escape the open cluster potential. Some neutron stars may be retained, but even in the more dense globular clusters this is only expected to be about 10%, so we might only expect a few percent in an open cluster. Of these that are retained, most will die as radio pulsars within a few Myr. There is a remote possibility that some of these pulsars will have very low magnetic fields (i.e. $\sim 10^{11}$), and might still be active today as their spin-down rates are very small. If one is still active, it still has to be beamed toward the Earth to be detectable. Slow pulsars tend to have small beaming fractions, so it is also quite unlikely. With a year of pulsar timing the position and period derivative should be measureable. If comparable to the age of the cluster (i.e., many Gyr) it would strengthen the case for the pulsar’s association. We therefore recommend a timing campaign be commenced on the pulsar.
Alternatively, the pulsar could come from a binary system, e.g. a low-mass X-ray binary, and was partially recycled before the binary was disrupted in the dynamic encounter with a cluster member (e.g. Verbunt & Freire, 2014). The scenario was used by Zhou et al. (2024) to explain the slow pulsars M15K and M15L in the core region of the globular cluster M15 (Wu et al., 2024; Zhou et al., 2024), although Kremer et al. (2024) argued that star clusters are not dense enough to effectively disrupt such binaries before they are spun up to milliseconds.
In addition to disrupted binaries, electron-captured supernovae (ECS) are also a possible pathway to producing slow pulsars in star clusters (Ivanova et al., 2008; Kremer et al., 2024; Singleton et al., 2024). Such supernovae could be triggered by the collapse of massive white dwarfs after accretion or merger. Since ECS is usually less energetic than core-collapse supernovae, pulsars formed in this way may suffer smaller kicks thus retain in the cluster (Lynch et al., 2012). Note that for the merger of white dwarfs, the systems could come from isolated binaries, and the merger could happen even in the case of low encounter rate.
Given the limited number of slow pulsars in star clusters, it is difficult to distinguish the various formation scenarios for this species. The scarcity of such pulsars could partly be attributed to a bias caused by search methods; the frequently adopted FFT method is less sensitive to long-period pulsars than to millisecond ones (e.g. Lazarus et al., 2015). Adopting a more advanced search method, e.g. the fast folding algorithm (Morello et al., 2020; Zhou et al., 2024), will likely discover more pulsars, including the slow ones, in both star clusters and test different formation mechanisms. We also provide a simple estimate of the expected number of recycled pulsars that could be hosted by NGC 6791 in Appendix A.
6 Conclusions
Pulsars located or born in open clusters have not drawn much attention in pulsar research, but may hold important clues in understanding the formation and evolution of different species of pulsars and star clusters. In this paper, we presented initial results from 20 hours of pulsar search observations in seven open clusters using FAST. We discovered a new pulsar J1922+37 in the direction of open cluster NGC 6791 (Fig. 2); it has a period of 1.92 second and a DM of $\sim 85$ pc cm-3. The NE2001 model gives a distance estimate of 4.79 kpc to the pulsar, consistent with the cluster distance, which ranges from 4.02 to 4.93 kpc in the literature. The pulsar is thus a plausible member of NGC 6791. If confirmed, PSR J1922+37 will be the first pulsar discovered in an open cluster.
Measurements of the pulsar proper motion will be crucial in determining whether PSR J1922+37 is associated with NGC 6791 or not but will be difficult. Assuming that old and massive open clusters such as NGC 6791 are similar to the Galactic globular clusters as pulsar birthplaces (e.g., the size of pulsar population is proportional to the stellar encounter rate), we estimate that around 9 recycled pulsars could be discovered in NGC 6791. We also highlight 10 open clusters that are promising sites for finding new pulsars (in Appendix B).
Follow-up observations have been planned with FAST, which could allow measurements of spin and astrometric parameters of PSR J1922+37, HI absorption, as well as a deeper search for more pulsars in NGC 6791. More pulsar discoveries with a similar DM to PSR J1922+37 in the region would provide compelling evidence that there is a population of detectable pulsars in open clusters. If so, their characteristic ages could be compared with that of the open cluster to learn how realistic pulsar characteristic ages are, and how long slow pulsars can remain detectable as radio pulsars.
We thank Lei Zhang, Pei Wang, Zhichen Pan, Lei Qian, Zu-Cheng Chen and Zhi-Qiang You for helpful discussions. This work made use of the data from FAST (Five-hundred-meter Aperture Spherical radio Telescope) (https://cstr.cn/31116.02.FAST). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. The data processing was mainly performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government. R.S. was supported by the National Science Foundation (NSF) grant AST-1816904. R.P.E. is supported by the Chinese Academy of Sciences President’s International Fellowship Initiative, grant No. 2021FSM0004. JPY, NW, WWZ are supported by the National Natural Science Foundation of China (grant No. 12041304, 12041303). WWZ is supported by National SKA Program of China (grant No. 2020SKA0120200). XJZ is supported by the National Natural Science Foundation of China (Grant No. 12203004) and by the Fundamental Research Funds for the Central Universities.
Appendix A Expected number of recycled pulsars in NGC 6791
@@
Following Bahramian et al. (2013), we estimate the expected number of recycled pulsar discoveries, $\lambda$, in NGC 6791 in three steps. First, using the cluster parameters in Section 5.1, we computed the central luminosity density, $\rho_{\rm c}$, according to the prescription given by Djorgovski (1993). Second, we obtained the encounter rate $\Gamma\sim\rho_{\rm c}^{2}r_{\rm c}^{3}/\sigma_{v}\approx 162$ assuming a King potential model (Dalessandro et al., 2015) and a velocity dispersion of $\sigma_{v}\sim 2$ km s-1. Note that $\Gamma$ has been normalized so that the encounter rate of 47 Tuc is $\Gamma_{\rm 47~{}Tuc}=1000$ (Bahramian et al., 2013). Finally, the expected number of pulsars is found by $\ln\lambda=-1.1+1.5\log_{10}\Gamma$ (Turk & Lorimer, 2013; Gao et al., 2024), which gives $\lambda\sim 9$ potential discoveries. This result is also in general agreement with the estimation using escape velocity; see Yin et al. (2024, Fig. 2).
Appendix B Potential pulsar discoveries in other open clusters
@@
The possible association between PSR J1922+37 and NGC 6791 motivates us to check if there are other open clusters that may also host pulsars. A detailed analysis inspecting the association between all known open clusters and pulsars is beyond the scope of this work, and will be presented in Zhou et al. (2024, in preparation). Here, we check the latest Galactic cluster catalog (Hunt & Reffert, 2024) and highlight the open clusters that are worthy of more careful pulsar surveys.
@@
We searched the catalog and picked out the open clusters that are similar to NGC 6791 in both total mass ($>10^{4}$ M) and core size ($<10$ pc), selecting open clusters with comparable core densities and thus encounter rate (e.g. Bahramian et al., 2013). We also set a lower age limit of 100 Myr, which is a typical lifetime for canonical pulsars.
@@
Ten open clusters, including NGC 6791, meet the above conditions and are shown in Table 4, which lists the total mass, angular size, and number of member stars. All the clusters are relatively close ($d<5$ kpc) and have a small core size with a typical value of $\sim 5\arcmin$. They also have a relatively high stellar concentration, with half of members within the tidal radius. These clusters are thus good candidates for future pulsar searches in open clusters.
@@
Table 4: Open clusters similar to NGC 6791 in order of decreasing total mass. Columns are cluster name, right ascension and declination, distance, age, total mass, the standard error of radial velocity, core radius (in pc and in arcmin), tidal radius, total radius, number of member stars within $r_{\rm tidal}$ and the total number of stars. The parameters of NGC 6791 are also included for comparison.